OpenStudy (anonymous):

what is the solution to 4^2x=about 3?

OpenStudy (anonymous):

x = –0.631 x = –0.396 x = 0.396 x = 0.631

OpenStudy (anonymous):

\[4^{2x}\approx3\]

OpenStudy (amistre64):

same as 8^x abt 3

OpenStudy (amistre64):

ln8/ln3

OpenStudy (anonymous):

so i just divide 8by 3 correct?

OpenStudy (amistre64):

or .. ln3/2 ln4 maybe more practical here

OpenStudy (amistre64):

4^2x = 3 ln(4^2x) = ln 3 2x ln(4) = ln 3 x = ln (3)/2 ln(4)

OpenStudy (anonymous):

ok b\c when i divideed 8 by 3 i got 2.66666666666666.7

OpenStudy (amistre64):

lol 4^2 = 16 not 8 .... forgot how to multiply

OpenStudy (anonymous):

opps i feel stupid now

OpenStudy (amistre64):

ln3/ln16 is good

OpenStudy (anonymous):

ok so it would be 0.1875

OpenStudy (amistre64):

your not working the ln function ...

OpenStudy (amistre64):

\[\frac{\ln3}{\ln16}\ne\frac{3}{16}\]

OpenStudy (anonymous):

i am completely lost

OpenStudy (amistre64):

how so, the answer is right there .... do you not know how to input it into a calcuator?

OpenStudy (anonymous):

no i dont have a caculator

OpenStudy (amistre64):

then thats prolly why you are not able to work the logarithm fucntion correctly

OpenStudy (anonymous):

thanks so much

OpenStudy (amistre64):

youre welcome; i spose we could have done some approximation techniques .... but they can get hairy

OpenStudy (anonymous):

lol i apppreciate the help my teacher get mad if i ask her for help